MICROWAVE SPECTRUM OF ^{12}C^{16}O

\underline{S.A. TASHKUN}, S.N. MIKHAILENKO, Laboratory of Theoretical Spectroscopy, V.E. Zuev Institute of Atmospheric Optics, 634055, Tomsk, RUSSIA.

12C^{16}O microwave spectrum of the 0-0,1-1,2-2, and 3-3 bands up to J=50 will be presented. The spectrum was calculated from an experimental dataset of energy levels^{a}. Calculated frequencies are given together with 99% confidence intervals. Comparison of the spectrum with microwave data containing in spectroscopic databanks will be discussed. This work was supported by CRDF (USA) Grant RUG1-2954-TO-09 and by RFBR. Grant 09-05-92508.

aS.A. Tashkun, T.I. Velichko, S.N. Mikhailenko, JQSRT, in press (2010).